6. 2.43 Consider the following hypothesis test 7.2.43 The following results are for two independent samples taken from two populations. Excel File: data10-03.xlsx Sample 1 Sample 2 n, 80 104 o 8.4 n 70 2 106 7.6 Enter negative values as negative numbers a. What is the value of the test statistic? (to 2 decimals) 1.53 b. What is the p-value? (to 4 decimals) 0630 c. With a-0.05, what is your hypothesis testing conclusion? Do not reject the null hypothesis v Hide Feedback Partially Correct Type here to search
Answer #1

The null and the alternative hypotheses are,

H_0: mu_1 - mu_2 = 0

H_a: mu_1 - mu_2 eq 0

a) Test statistic is, z = -1.53

b) p-value = 2 * P(Z < -1.53) = 2 * 0.0630 = 0.1260

p-value = 0.1260

c) Here, p-value = 0.1260 > a-0.05

So, we do not reject the null hypothesis ( H0).